# HSACNet: Hierarchical Scale-Aware Consistency Regularized Semi-Supervised Change Detection

### v1

@@ Abstract
Index Terms:
Change Detection (CD), Semi-Supervised Learning, Segment Anything Model, Scale-Aware
@@ III-B Change Detection Network
SAM2 Image Encoder with Adapters. Given the powerful feature extraction capability of SAM2, we select its Hiera-T [9] backbone with 28M parameters as the basis for our siamese encoder. It has 4 layers with channels $=[96,192,384,768]$ and blocks $=[1,2,7,2]$. This hierarchical structure excels at extracting inter-layer multi-scale features from input image pairs. The process is expressed as:
@@
Here, $\text{CBR}_{k,g}(\cdot)$ denotes a convolution layer with kernel $k\times k$ and group $g$ (default $g=1$), followed with Batch Normalization and ReLU, where $k=2j-1$, $g=2^{j}$. The symbol ||$\cdot$|| ⋅ | indicates the absolute operation.
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the channel-wise concatenation. The refined change feature $\textbf{D}^{SADAM}$ is obtained by:
@@ III-C Our Proposed HSACNet
$IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$
@@
$IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$ $IoU^{c}$ $OA$
AdvEnt [12] 56.7 95.52 57.5 95.99 70.3 97.28 70.8 97.29 53.2 95.62 56.5 96.26 59.4 96.62 64.1 96.83
@@ IV-A Experimental Setups
Datasets. We conduct experiments on four benchmark remote sensing CD datasets, namely WHU-CD [17], LEVIR-CD [18], GZ-CD [5], and EGY-BCD [19]. WHU-CD [17] has two VHR aerial image pairs (2012 and 2016, Christchurch, New Zealand), with $32507\times 15354$ pixels and 0.075 m pixel resolution. LEVIR-CD contains 637 VHR image pairs ($1024\times 1024$ pixels, 0.5 m resolution), showing land use changes (5-14 years, Texas, USA). GZ-CD offers 19 high-resolution, season-varying image pairs (2006-2019, Guangzhou suburbs), sized $1006\times 1168$ to $4936\times 5224$ pixels and 0.55 m resolution. EGY-BCD comprises 6091 VHR image pairs (2017-2022, four Egyptian regions), $256\times 256$ pixels and 0.25 m resolution. All images are cropped into $256\times 256$ non-overlapping patches and split into training, validation, and test sets. The training set is further divided into labeled and unlabeled data at ratios of $[5\%,95\%]$, $[10\%,90\%]$, $[20\%,80\%]$, $[40\%,60\%]$.
@@
Metrics. We adopt intersection over union ($IoU^{c}$) and overall accuracy ($OA$) to evaluate performance. $IoU^{c}$ quantifies precision by measuring the overlap between predicted and actual changes, while $OA$ offers a general accuracy measure. Their formulas are $IoU^{c}=\frac{TP}{TP+FP+FN}$, $OA=\frac{TP+TN}{TP+FP+FN+TN}$, with $TP$, $TN$, $FP$ and $FN$ signifying true positive, true negative, false positive, and false negative.
Figure 5: Detection results of different methods on four CD datasets at $5\%$ labeled ratio. Here, white for $TP$, black for $TN$, red for $FP$, and green for $FN$.
@@ IV-C Ablation Studies
Method $IoU^{c}$ $OA$ GFLOPs Params (M)

### v2

@@ III-B Change Detection Network
SAM2 Image Encoder with Adapters. Given the powerful feature extraction capability of SAM2, we select its Hiera-T [9] backbone with 28M parameters as the basis for our siamese encoder. It has 4 layers with channels =[96,192,384,768]=[96,192,384,768] and blocks =[1,2,7,2]=[1,2,7,2]. This hierarchical structure excels at extracting inter-layer multi-scale features from input image pairs. The process is expressed as:
@@
Here, $\text{CBR}_{k,g}(\cdot)$ denotes a convolution layer with kernel $k\times k$ and group $g$ (default $g=1$), followed with Batch Normalization and ReLU, where k=2j1k=2j-1, $g=2^{j}$. The symbol $|\cdot|$ indicates the absolute operation.
@@
where $\cdot$$\cdot$] denotes the channel-wise concatenation. The refined change feature $\textbf{D}^{SADAM}$ is obtained by:
@@ III-C Our Proposed HSACNet
$IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA
@@
$IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA $IoU^{c}$ OAOA
AdvEnt [12] 56.7 95.52 57.5 95.99 70.3 97.28 70.8 97.29 52.0 95.30 58.1 96.26 59.8 96.46 63.8 96.94
@@ IV-A Experimental Setups
Datasets. We conduct experiments on four benchmark remote sensing CD datasets, namely WHU-CD [17], LEVIR-CD [18], GZ-CD [5], and EGY-BCD [19]. WHU-CD [17] has two VHR aerial image pairs (2012 and 2016, Christchurch, New Zealand), with $32507\times 15354$ pixels and 0.075 m pixel resolution. LEVIR-CD contains 637 VHR image pairs ($1024\times 1024$ pixels, 0.5 m resolution), showing land use changes (5-14 years, Texas, USA). GZ-CD offers 19 high-resolution, season-varying image pairs (2006-2019, Guangzhou suburbs), sized $1006\times 1168$ to $4936\times 5224$ pixels and 0.55 m resolution. EGY-BCD comprises 6091 VHR image pairs (2017-2022, four Egyptian regions), $256\times 256$ pixels and 0.25 m resolution. All images are cropped into $256\times 256$ non-overlapping patches and split into training, validation, and test sets. The training set is further divided into labeled and unlabeled data at ratios of $5\%,95\%]$, $10\%,90\%]$, $20\%,80\%]$, $40\%,60\%]$.
@@
Metrics. We adopt intersection over union ($IoU^{c}$) and overall accuracy ($OA$) to evaluate performance. $IoU^{c}$ quantifies precision by measuring the overlap between predicted and actual changes, while $OA$ offers a general accuracy measure. Their formulas are $IoU^{c}=\frac{TP}{TP+FP+FN}$, $OA=\frac{TP+TN}{TP+FP+FN+TN}$, with $TP$, TNTN, FPFP and $FN$ signifying true positive, true negative, false positive, and false negative.
Figure 5: Detection results of different methods on four CD datasets at $5\%$ labeled ratio. Here, white for TPTP, black for TNTN, red for FPFP, and green for FNFN.
@@ IV-C Ablation Studies
Method $IoU^{c}$ OAOA GFLOPs Params (M)